A car is traveling at a speed of 40.6 m/s on an interstate highway where the speed limit is 65.0 mi/h. Is the driver exceeding the speed limit?

& then justify your answer.
Answer #1

Given that

speed of car v =40.6 m/s

we know that

1 hour = 3600 seconds

so in one hour car travelling = 40.6*3600 = 146160 meter

we know that 1 miles = 1609.34 meter

so in one hour car will travel = 146160/1609.34= 90.81 miles

so it clearly shows that car exceeding the limit of speed

Good luck